Show by the method of variation of parameters that the general solution of the differential equation -y''=f(x) can be written in the form y = φ ( x) = c1+c2x-∫0x(x-s)f(s)ds where c1 and c2 are arbitrary constants.


1
Expert's answer
2021-08-09T16:24:56-0400

"\displaystyle
y" = -f(x)\\\\\ny" = 0\\
y' = C_2\\
y = C_2 + C_1x\\

\textsf{Wronskian}(1, x) = 1\\

C_2 = -\int \frac{x \times -f(x)}{1} \mathrm{d}x = \int xf(x)\,\,\mathrm{d}x + c_2\\
 
C_1 = \int \frac{1 \times -f(x)}{1} \mathrm{d}x = -\int f(x)\,\,\mathrm{d}x + c_1\\

\begin{aligned}
y &= -x\left(\int f(x)\,\,\mathrm{d}x + c_1\right) + \int xf(x)\,\,\mathrm{d}x + c_2
\\&= c_1x + c_2 - x\left(\int_0^x f(s)\,\,\mathrm{d}s\right) + \int_0^x sf(s)\,\,\mathrm{d}s
\\&= c_1x + c_2 - \int_0^x xf(s)\,\,\mathrm{d}s + \int_0^x sf(s)\,\,\mathrm{d}s
\\&= c_1x + c_2 - \int_0^x (x - s)f(s)\,\,\mathrm{d}s
\end{aligned}"